A right triangle has one 37 degree angle. Find the other two angles.
Answer by Theo(3464)

sum of the angles of any triangle is 180 degrees
right triangle has one angle of 90 degrees
90 + 37 = 127
180 - 127 = 53
the other angle is 53 degrees.
90 + 37 + 53 = 180